OpenStudy (anonymous):

Evaluate the following definite integral using the Fundamental Theorem on Calculus: \[\int\limits_{1}^{4}1divx ^{2}dx\]

OpenStudy (amistre64):

\div

OpenStudy (amistre64):

or simply: \frac{1}{x^2}

OpenStudy (amistre64):

it seems easier to just convert to exponents: 1/x^2 = x^-2

OpenStudy (jamesj):

Let \[ F(x) = \int \frac{1}{x^2} dx \] By the F.T.Calculus, \( dF/dx = 1/x^2 = x^{-2} \). Hence the question is: what function F has that property?

OpenStudy (jamesj):

So ... what's your answer to the first part here? What's the function F(x)?

OpenStudy (anonymous):

is that the antiderivative of f(x)?

OpenStudy (jamesj):

If f(x) = 1/x^2, then F(x) is indeed the anti-derivative.

OpenStudy (anonymous):

James, i don't think I can do this one :(

OpenStudy (jamesj):

What is the derivative of 1/x ?

OpenStudy (anonymous):

That is x^-1 (i think)

OpenStudy (jamesj):

yes. Hence it's derivative is what?

OpenStudy (anonymous):

1/x?

OpenStudy (jamesj):

If g(x) = 1/x = x^(-1), then dg/dx = ... what? Hint: what is the derivative of h(x) = x^n, for an arbitrary number n?

OpenStudy (jamesj):

If h(x) = x^n, then dh/dx = nx^(n-1). Therefore what is dg/dx?

OpenStudy (anonymous):

is it -x^-2

OpenStudy (jamesj):

Correct. Now we want a function F(x) such that dF/dx = +x^(-2). What then must that function be?

OpenStudy (anonymous):

the function must be: -1/3 x^-3??

OpenStudy (jamesj):

No. You know that (d/dx) 1/x = -1/x^2. You want a function F so that (d/dx) F = +1/x^2. But +1/x^2 = (-1)(-1/x^2). Therefore F must equal ... what?

OpenStudy (jamesj):

Ok. let's try this one other way. What is the derivative of g(x) = A/x, where A is a real number, a constant?

OpenStudy (jamesj):

dg/dx = ...what?

OpenStudy (anonymous):

that is Ax^-1?

OpenStudy (jamesj):

yes.

OpenStudy (jamesj):

Quickly now. You've already done this for x^(-1)

OpenStudy (jamesj):

If g(x) = A/x = Ax^(-1), then dg/dx = -A/x^2 = -Ax^(-2)

OpenStudy (anonymous):

dg/gx= -Ax^-2

OpenStudy (jamesj):

That being the case, for what value of A is dg/dx = 1/x^2 ?

OpenStudy (anonymous):

A=1?

OpenStudy (jamesj):

No. If A = 1, then dg/dx = -1/x^2

OpenStudy (anonymous):

A= -1

OpenStudy (jamesj):

Yes. That being the case, the anti-derivative of 1/x^2 is what?

OpenStudy (jamesj):

The anti-derivative is the function F such that dF/dx = 1/x^2

OpenStudy (anonymous):

the anti-derivative of 1/x^2=> F(x)=1/2x^3??

OpenStudy (anonymous):

i think I just dis all wrong

OpenStudy (jamesj):

No. dF/dx = x^(-2). What is the derivative of 1/x?

OpenStudy (anonymous):

x^-1

OpenStudy (jamesj):

1/x = x^(-1). What is the derivative of that function wrt x?

OpenStudy (anonymous):

it is -x^-2?

OpenStudy (jamesj):

(d/dx)(1/x) = -1/x^2 = -x^(-2). Because writing 1/x = x^(-1), it's derivative is (-1).x^(-1-1) = -x^(-2)

OpenStudy (jamesj):

So what again is the derivative of g(x) = A/x ?

OpenStudy (anonymous):

Ax^-1

OpenStudy (jamesj):

that is NOT the derivative.

OpenStudy (jamesj):

If g(x) = A/x dg/dx = ...

OpenStudy (anonymous):

oops... -Ax^-2?

OpenStudy (jamesj):

dg/dx = -A/x^2, yes. Hence for what value of A is dg/dx = 1/x^2 ?

OpenStudy (anonymous):

A= -1

OpenStudy (jamesj):

Correct. Hence what is the function F such that dF/dx = 1/x^2 ?

OpenStudy (anonymous):

F=1?

OpenStudy (jamesj):

If F = 1, dF/dx = 0. With our A = -1, the function g is g(x) = -1/x. What is the derivative dg/dx in this case?

OpenStudy (anonymous):

the derivative of dg/dx=1^-2

OpenStudy (jamesj):

1x^(-2) = 1/x^2 I.e., if g(x) = -1/x, then dg/dx = 1/x^2. Hence what is the antiderivative function F that has the property dF/dx = 1/x^2 ?

OpenStudy (anonymous):

is it -1?

OpenStudy (jamesj):

Kewnisha: have a coffee or a glass of coke or something. Didn't we just show that if g(x) = -1/x, then dg/dx = 1/x^2? So mustn't a function F which has the property dF/dx = 1/x^2 be exactly the same?

OpenStudy (jamesj):

That is, F(x) = -1/x.

OpenStudy (anonymous):

OHHH... now I get it... ouch... that was really bad :(

OpenStudy (jamesj):

Ok. Returning now to your integral, \[ \int_1^4 \frac{1}{x^2} dx = F(4) - F(1) \] where dF/dx = 1/x^2

OpenStudy (jamesj):

Therefore the integral is equal to what? F(4) - F(1) = ...

OpenStudy (anonymous):

-1/4 - -1/1 => -1/4+1= 3/4?

OpenStudy (jamesj):

Correct.

OpenStudy (jamesj):

So tell me again, what is the derivative of 1/x?

OpenStudy (anonymous):

that is x^-1

OpenStudy (jamesj):

1/x = x^(-1), yes. What is its derivative?

OpenStudy (anonymous):

its derivative if -x^(-2)

OpenStudy (jamesj):

The derivative of 1/x is -1/x^2, yes. What is the derivative of -1/x?

OpenStudy (anonymous):

the derivative of -1/x is x^-2

OpenStudy (jamesj):

1/x^2 yes. What is the derivative of A/x ?

OpenStudy (jamesj):

..where A is a constant?

OpenStudy (anonymous):

Ax^-1

OpenStudy (jamesj):

the derivative of A/x is ...

OpenStudy (anonymous):

it is Ax^-1

OpenStudy (jamesj):

(d/dx)(A/x) = -A/x^2

OpenStudy (jamesj):

What's the derivative of 2/x?

OpenStudy (anonymous):

2x^-1

OpenStudy (jamesj):

No. What is (d/dx)(1/x)?

OpenStudy (jamesj):

(d/dx)(1/x) = (d/dx)(x^-1) = -x^-2 = -1/x^2

OpenStudy (jamesj):

Hence what is (d/dx)(2/x)?

OpenStudy (anonymous):

-2/x^2

OpenStudy (jamesj):

correct. What is (d/dx)(A/x) ?

OpenStudy (anonymous):

that is: -A/x^2

OpenStudy (jamesj):

Yes. What is the (d/dx)(-1/x) ?

OpenStudy (anonymous):

1/x^2

OpenStudy (jamesj):

yes. What is the anti-derivative of 1/x^2 ?

OpenStudy (anonymous):

-1/x

OpenStudy (jamesj):

Yes. What is the anti-derivative of -2/x^2 ?

OpenStudy (anonymous):

2/x

OpenStudy (jamesj):

Yes. What is (d/dx)(x^2) ?

OpenStudy (anonymous):

-2/x^3

OpenStudy (jamesj):

x^2, not 1/x^2

OpenStudy (anonymous):

ohh... 2x

OpenStudy (jamesj):

What's that antiderivative of 1/x^3 ?

OpenStudy (anonymous):

that is: -3x^-4

OpenStudy (jamesj):

No. What's (d/dx)(1/x^2) ?

OpenStudy (anonymous):

-1/x

OpenStudy (jamesj):

No, you wrote it down a moment ago by accident. (d/dx)(1/x^2) = ...

OpenStudy (anonymous):

-2/x^3

OpenStudy (jamesj):

Yes. Hence the anti-derivative of (-2/x^3) is what?

OpenStudy (anonymous):

1/x^2

OpenStudy (jamesj):

Yes, good. What then is the anti-derivative of 1/x^3 ?

OpenStudy (anonymous):

1/4 x^-4